OpenStudy (anonymous):

The fourteen students in Mateo's computer programming class are using the computer lab. Some computers are shared by pairs of students, and 2 students work alone. Which equation could you use to determine how many computers are shared in the computer lab? Let c be the number of shared computers. the quantity c over two all minus two equals fourteen two times c minus two equals fourteen two times c plus two equals fourteen the quantity c over two all plus two equals fourteen

OpenStudy (anonymous):

@Yttrium

ganeshie8 (ganeshie8):

can i give a hand ? :)

OpenStudy (anonymous):

Anyone can

ganeshie8 (ganeshie8):

:D say the number of computers shared = \(c\)

ganeshie8 (ganeshie8):

so that means, each of these will get 2 students. so number of students that are on shared computers = \(2 \times c\)

ganeshie8 (ganeshie8):

next, how many students are working alone ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

@AkashdeepDeb

ganeshie8 (ganeshie8):

correct, simply write equation for total number of students nw

ganeshie8 (ganeshie8):

total number of students = nuber of students sharing computers+ number of students working alone

ganeshie8 (ganeshie8):

= ?

OpenStudy (anonymous):

total number of students 14 number of students sharing computers 12 number of students working alone 2 14 = 12 + 2

ganeshie8 (ganeshie8):

you're correct, but dont simplify fullly..

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

il change wat u wrote above

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

total number of students \(14\), number of students sharing computers \(2 \times c\) number of students working alone \(2\) \(2 \times c + 2 = 14\)

OpenStudy (anonymous):

2 x 6 + 2 = 14

ganeshie8 (ganeshie8):

dont do that,

ganeshie8 (ganeshie8):

look at ur options

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so its 2c + 2 = 14 is the answer right

ganeshie8 (ganeshie8):

yes !